OpenStudy (anonymous):

Find the product. (n^3)^2 · (n^5)^4

OpenStudy (whpalmer4):

You need two properties of exponents: \[(x^a)^b = x^{a*b}\]and\[x^a*x^b = x^{a+b}\]

OpenStudy (anonymous):

I just need to figure out the total power of n. The answer is n to the _ power.

OpenStudy (anonymous):

n^3)^3=n^9 (multiply when an exponent has another exponent). thus, (n^4)^5=n^20 n^9*n^20=n^29 (add when multiplying exponents).

OpenStudy (anonymous):

Thank you :)